### v2

@@ 0.A.5 Formal abstraction of the learning-based controller
0.A.6 Utility Functions for Strategic Withholding
@@
0.A.7 Risk Metrics for Default and Liquidation
@@
$\pi_{t}(p_{t+1})\coloneqq\sum_{i\in\text{Borrowers}}\max\left\{0,B_{t}(i)-C_{t}(i)\cdot p_{t+1}\right\}$ (20)
@@
Appendix 0.B Complimentary Material of Evaluation
0.B.1 Estimated parameters from Aave demand and supply data
Figure 4: Estimated parameters for the demand curve. The top plot shows the estimated $\mathtt{a}_{t}^{b}$ parameter over time for each token, and the bottom plot shows the estimated $\mathtt{b}_{t}^{b}$ parameter over time for each token. RLS algorithm was used with forgetting factor $0.8$.
Figure 5: Estimated parameters for the supply curve. The top plot shows the estimated $\mathtt{a}_{t}^{l}$ parameter over time for each token, and the bottom plot shows the estimated $\mathtt{b}_{t}^{l}$ parameter over time for each token. RLS algorithm was used with forgetting factor $0.8$.
The statistics of the estimated parameters for demand and supply are presented respectively in Table 2 and Table 3. Contrary to expectations, the average of $\mathtt{a}^{b}$ is positive, meaning borrowers are more willing to borrow at higher rates. This can be attributed to the complex reward systems employed by current DeFi borrowing and lending protocols. These protocols often provide a portion of their native tokens to users as rewards and offer additional promotions. As a result, the net profit from borrowing can sometimes be positive, encouraging borrowers to borrow more, even at high interest rates. It also appears that, due to other factors influencing lenders’ choices, $\mathtt{a}^{l}$ is negative, contrary to expectations. In general, the overly complicated reward systems of current DeFi platforms lead to highly complex and difficult-to-understand user behavior. A simpler, more principled design for DeFi borrowing and lending, like the one we propose in this paper, could make user behavior much more interpretable.
Token a - Mean Absolute Change (Between Timeslots) a - Mean a - Variance
DAI 53797.3060 1506059.7038 1904821040691.7227
USDC 214256.9283 7477767.8699 16233722878803.0078
WETH 439991.9727 32806929.4749 372275477524692.6875
Token b - Mean Absolute Change (Between Timeslots) b - Mean b - Variance
DAI 133892.0375 9945763.0576 43141197173355.0391
USDC 513954.1723 63701880.4670 2261658367622412.0000
WETH 1175699.9110 79294641.5824 2347228562235640.5000
Table 2: Statistics of demand parameters from 2/2023 to 7/2023
Token a - Mean Absolute Change (Between Timeslots) a - Mean a - Variance
DAI 217286.5234 -2087579.7981 9953611878125.8164
USDC 656437.1639 -6774126.9809 46222689182475.4531
WETH 1405347.$3$10153351.0180 511814384630384.3125
Token b - Mean Absolute Change (Between Timeslots) b - Mean b - Variance
DAI 328886.3333 26011504.3014 174238744066342.8750
USDC 1002216.0937 129651988.7290 4685211195900898.0000
WETH 2302670.3581 362200857.1465 37906774777703448.0000
Table 3: Statistics of supply parameters from 2/2023 to 7/2023
0.B.2 Utilization optimization
Figure 7 illustrates how the RLS-based controller operates on user behavior parameters learned from the actual data of Aave’s DAI pool. The blue line represents the realized utilization in the protocol, while the green line indicates the desired utilization set by the protocol. It is evident that Aave’s protocol struggles to effectively stabilize at the desired utilization, whereas the RLS-based algorithm demonstrates superior stability in achieving this goal.
@@ 0.B.2 Utilization optimization
Figure 6: Impact of forgetting factor on the mean square error of the estimated parameters by the RLS-based algorithms.
Figure 7: Comparison of utilization between RLS-based controller and Aave’s static curves, with user supply and demand curves learned from real Aave DAI pool data, $\rho=0.8$.
0.B.3 Supply maximizing
@@
Figure 8: Comparison of supply, used as a proxy for revenue, between the RLS-based controller and Aave’s static curves. The parameters evolve according to a random walk with Gaussian noise. The x-axis represents the relative standard deviation of the noise in percentage. The fixed factor is set to $\rho=0.8$.
Appendix 0.C Proofs
0.C.1 Proof of theorem 3.1
Proof
Our proof is applying a slightly modified version of the proof provided in [14]. The condition required for convergence of the RLS algorithm is that the input vector of the linear system should be persistently excited. In other words if we want to estimate $\mathbf{\theta}_{t}$, and the linear system is $y_{t}=\mathbf{x}_{t}^{T}\mathbf{\theta}_{t}+e_{t}$ then $\mathbf{x}_{t}$ should hold in:
@@ Appendix 0.C Proofs
The above analysis shows that as long as for any $t$, there is some $N$ such that Algorithm 2 sets at least two distinct interest rates from time $t$ to $t+N$, then the excitation condition of paper [14] is met.
@@
So now we can apply the results of [14] for the convergence of the RLS algorithm to the right $\theta_{t}$. They write the estimation error $\tilde{\theta}_{t}=\theta_{t}-\hat{\theta}_{t}$ as sum of three terms $\tilde{\theta}_{t}\coloneqq\tilde{\theta}^{1}_{t}+\tilde{\theta}^{2}_{t}+\tilde{\theta}^{3}_{t}$. Given that $\theta$ has some major change at time $t_{0}$, the question is at time $t>t_{0}$ what is an upper bound of each of $\tilde{\theta}^{1}_{t},\tilde{\theta}^{2}_{t},\tilde{\theta}^{3}_{t}$? They show that the first and second terms decay exponentially over time i.e., $\|\tilde{\theta}^{1}_{t}+\tilde{\theta}^{2}_{t}\|=\mathcal{O}(\rho^{t})$ whereas the third term $\tilde{\theta}_{t}^{3}$ which is the only term that is a function of the observation noise, is as follows:
$\limsup_{t\to\infty}\mathbb{E}\left(\|\tilde{\theta}_{t}^{3}\|^{2}\right)\leq\nu^{2}\frac{M}{m^{2}}\left(\frac{1}{\rho^{N}}-1\right)^{2}\frac{1}{\ln\left(\frac{1}{\rho}\right)}=\mathcal{O}\left(\frac{\nu^{2}}{\rho^{2N}\ln{\frac{1}{\rho}}}\right)$
@@
0.C.2 Proof of theorem 3.2
Proof
@@
$U$ $=\frac{-\mathtt{a}^{b}\left(R_{\text{slope1}}+R_{\text{slope2}}\left(\frac{U-U^{*}}{1-U^{*}}\right)\right)+\mathtt{b}^{b}}{L}\implies U=\frac{\mathtt{b}^{b}-\mathtt{a}^{b}R_{\text{slope1}}+\frac{\mathtt{a}^{b}R_{\text{slope2}}U^{*}}{1-U^{*}}}{L+\frac{\mathtt{a}^{b}R_{\text{slope2}}}{1-U^{*}}}$ (21)
$\implies r=R_{\text{slope1}}+\frac{R_{\text{slope2}}\left(\mathtt{b}^{b}-\mathtt{a}^{b}R_{\text{slope1}}-U^{*}L\right)}{\left(L+\frac{\mathtt{a}^{b}R_{\text{slope2}}}{1-U^{*}}\right)(1-U^{*})}$ (22)
@@
$\mathcal{R}_{t}^{\mathcal{P},\nu}(U^{*};\{\theta,\omega\})=\begin{cases}\left|R_{\text{slope1}}+\frac{R_{\text{slope2}}\left(\mathtt{b}^{b}-\mathtt{a}^{b}R_{\text{slope1}}-U^{*}L\right)}{\left(L+\frac{\mathtt{a}^{b}R_{\text{slope2}}}{1-U^{*}}\right)(1-U^{*})}-\frac{\mathtt{b}^{b}-L\,U^{*}}{\mathtt{a}^{b}}\right|&\text{if }R_{\text{slope1}}<\frac{\mathtt{b}^{b}-L\,U^{*}}{\mathtt{a}^{b}},\\ \\ \left|\frac{\mathtt{b}^{b}R_{\text{slope1}}}{U^{*}\,L+\mathtt{a}^{b}R_{\text{slope1}}}-\frac{\mathtt{b}^{b}-L\,U^{*}}{\mathtt{a}^{b}}\right|&\text{if }R_{\text{slope1}}>\frac{\mathtt{b}^{b}-L\,U^{*}}{\mathtt{a}^{b}}.\end{cases}$
$\mathcal{R}_{t}^{\mathcal{P},\nu}(U^{*};\{\theta,\omega\})=\begin{cases}\frac{\Delta}{1+\frac{\mathtt{a}^{b}\cdot R_{\text{slope1}}}{L\,U^{*}}}&\text{if }R_{\text{slope1}}<\frac{\mathtt{b}^{b}-L\,U^{*}}{\mathtt{a}^{b}},\\ \\ \frac{\Delta}{1+\frac{\mathtt{a}^{b}\cdot R_{\text{slope2}}}{L(1-U^{*})}}&\text{if }R_{\text{slope1}}>\frac{\mathtt{b}^{b}-L\,U^{*}}{\mathtt{a}^{b}}.\end{cases}$
$\mathcal{R}_{t}^{\mathcal{P},\nu}(U^{*};\{\theta,\omega\})\geq\frac{\Delta}{1+\max\{\frac{\mathtt{a}^{b}}{L}\}\cdot\max\{\frac{R_{\text{slope1}}}{U^{*}},\frac{R_{\text{slope2}}}{1-U^{*}}\}}$
0.C.3 Proof of theorem 0.A.1
Proof
@@
$U_{t}=\frac{\mathtt{b}^{l}+\sqrt{(\mathtt{b}^{l})^{2}-4\mathtt{a}^{l}r_{t}(\mathtt{a}^{b}r_{t}-\mathtt{b}^{b})}}{2\mathtt{a}^{l}r_{t}}$ (23)
@@
$r=\frac{\mathtt{b}^{b}+\sqrt{(\mathtt{b}^{b})^{2}+4(\mathtt{a}^{b})^{2}(\mathtt{b}^{l})^{2}}}{2\mathtt{a}^{b}}$ (24)
@@
0.C.4 Proof of Proposition 1
Proof
@@ 0.C.4 Proof of Proposition 1
0.C.5 Proof of Theorem 4.1
Proof
@@
$\text{Utility}_{b}^{\text{honest}}\geq$ $-\Bigg{(}\delta_{b}\mathtt{b}^{b}\tau\zeta\frac{r^{b}-r_{\text{min}}}{r_{\text{max}}-r_{\text{min}}}\frac{r^{b}+r_{\text{min}}}{2}+C_{\tau,\zeta}^{r^{b}}+\delta_{b}\mathtt{b}^{b}(T-\tau)r^{b}\Bigg{)}$ (25)
If the borrower follows the honest strategy and reports $r^{b}$, during the exploration phase, they will borrow whenever $r\in[r_{\text{min}},r^{b}]$ and pay an average interest rate of $\frac{r_{\text{min}}+r^{b}}{2}$ during those timeslots. There are $\tau\zeta\frac{r^{b}-r_{\text{min}}}{r_{\text{max}}-r_{\text{min}}}$ such timeslots based on the uniform randomization of $\mathcal{C}_{D}$ described in Appendix [0.A.5]. During the rest of the exploration phase, i.e., when $r>r^{b}$ or during the non-randomized times, we denote the cost of the borrower by $C_{\tau,\zeta}^{r^{b}}$. Finally, during the exploitation phase $T-\tau$, the borrower pays at most rate $r^{b}$ (potentially even lower depending on the rest of the parameters and objective utilization $U^{*}$).
@@
$+C_{\tau,\zeta}^{r^{b}}+\delta_{b}\mathtt{b}^{b}(T-\tau)\bar{r}^{b}\Bigg{)}$ (26)
@@
$\text{Utility}_{b}^{\text{adv}}-\text{Utility}_{b}^{\text{honest}}\leq$ $-\delta_{b}\mathtt{b}^{b}\Bigg{(}\frac{\tau\zeta}{2}\frac{(r^{b}-\bar{r}^{b})^{2}}{r_{\text{max}}-r_{\text{min}}}+\big{(}T-\tau\big{)}(\bar{r}^{b}-r^{b})\Bigg{)}$ (27)
$\text{Utility}_{b}^{\text{adv}}-\text{Utility}_{b}^{\text{honest}}\geq 0$ (28)
$\implies\Bigg{(}\frac{\tau\zeta}{2}\frac{(r^{b}-\bar{r}^{b})^{2}}{r_{\text{max}}-r_{\text{min}}}+\big{(}T-\tau\big{)}(\bar{r}^{b}-r^{b})\Bigg{)}\leq 0$ (29)
$\implies 0<r^{b}-\bar{r}^{b}<\frac{2(r_{\text{max}}-r_{\text{min}})}{\zeta}\left(\frac{T}{\tau}-1\right)$ (30)
@@
    $\frac{\mathtt{b}^{b}(1-\delta_{b})-\mathtt{a}^{b}\,\bar{r}}{L}<U^{*}\implies\bar{r}>\frac{\mathtt{b}^{b}(1-\delta_{b})-L\,U^{*}}{\mathtt{a}^{b}}$ (31)
@@
    $U^{*}-\frac{\mathtt{b}^{b}(1-\delta_{b})-\mathtt{a}^{b}\,\bar{r}}{L}>\frac{\mathtt{b}^{b}-\mathtt{a}^{b}\,\bar{r}}{L}-U^{*}\implies\bar{r}>\frac{\mathtt{b}^{b}(1-\frac{\delta_{b}}{2})-L\,U^{*}}{\mathtt{a}^{b}}$ (32)
@@
$\text{Utility}_{l}^{\text{honest}}\geq$ $\Bigg{(}\delta_{l}L\tau\zeta\frac{r_{\text{max}}-r^{l}}{r_{\text{max}}-r_{\text{min}}}\frac{r^{l}+r_{\text{max}}}{2}+C_{\tau,\zeta}^{r^{l}}+\delta_{l}L(T-\tau)r^{l}\Bigg{)}$ (33)
@@
$+C_{\tau,\zeta}^{r^{l}}+\delta_{l}L(T-\tau)\bar{r}^{l}\Bigg{)}$ (34)
@@
    $\frac{\mathtt{b}^{b}-\mathtt{a}^{b}\,\bar{r}}{L(1-\delta_{l})}>U^{*}\implies\bar{r}<\frac{\mathtt{b}^{b}-L(1-\delta_{l})\,U^{*}}{\mathtt{a}^{b}}$ (35)
@@
    $\frac{\mathtt{b}^{b}-\mathtt{a}^{b}\,\bar{r}}{L(1-\delta_{l})}-U^{*}>U^{*}-\frac{\mathtt{b}^{b}-\mathtt{a}^{b}\,\bar{r}}{L}\implies\bar{r}<\frac{\mathtt{b}^{b}-U^{*}\,L\frac{(1-\delta_{l})}{(1-\frac{\delta_{l}}{2})}}{\mathtt{a}^{b}}$ (36)
@@
0.C.6 Proof of Theorem 4.2
Proof
@@
$\hat{L}_{\text{strtgc}}>0\ \implies r_{o}^{l}\leq\frac{B\alpha}{L(1-\delta_{l})}+\beta$ (37)

### v3

@@ A.3. Formal abstraction of the learning-based controller
A.4. Utility Functions for Strategic Withholding
@@
A.5. Risk Metrics for Default and Liquidation
@@
(21) $\pi_{t}(p_{t+1})\coloneqq\sum_{i\in\text{Borrowers}}\max\left\{0,B_{t}(i)-C_{t}(i)\cdot p_{t+1}\right\}$
@@
A.6. M-estimate function for robust RLS
We adopt the M-estimate function proposed by (Zou et al., 2000):
(22) $\phi(e)\triangleq\begin{cases}\frac{e^{2}}{2},&0<|e|<\xi,\\[8.0pt] \xi|e|-\frac{\xi^{2}}{2},&\xi\leq|e|<\Delta_{1},\\[8.0pt] \frac{\xi}{2}(\Delta_{2}+\Delta_{1})-\frac{\xi^{2}}{2}+\frac{\xi(|e|-\Delta_{2})^{2}}{\Delta_{1}-\Delta_{2}},&\Delta_{1}\leq|e|<\Delta_{2},\\[8.0pt] \frac{\xi}{2}(\Delta_{2}+\Delta_{1})-\frac{\xi^{2}}{2},&\Delta_{2}\leq|e|.\end{cases}$
Following (Zou et al., 2000), the thresholds $\xi$, $\Delta_{1}$, and $\Delta_{2}$ are determined using confidence levels of $0.050.025$, and $0.01$. Under the assumption of Gaussian noise, these thresholds are computed as $\xi=1.96\hat{\sigma}(t)$, $\Delta_{1}=2.24\hat{\sigma}(t)$, and $\Delta_{2}=2.576\hat{\sigma}(t)$, where $\hat{\sigma}(t)$ is the estimated standard deviation of the noise.
Appendix B Complimentary Material of Evaluation
B.1. Parameter estimation of Compound demand and supply curves
We analyze data from the three major lending pools—DAI, USDC, and USDT—on Compound V2 over the period from February 9, 2024, to February 9, 2025. Compound V2 uses different interest rates for borrowing and supplying assets. To model the relationship between past interest rates and subsequent demand and supply, we estimate the functional dependence between the interest rate at time $t-\Delta$ and the corresponding demand and supply at time $t$.
A key challenge is determining the appropriate time lag $\Delta$, as the response time of users to interest rate changes is not known a priori. To identify a reasonable lag, we evaluate different values of $\Delta$ ranging from 1 to 30 time slots (each time slot corresponding to a three-hour interval). For each candidate $\Delta$, we apply Recursive Least Squares (RLS) with a forgetting factor of $\rho=0.95$ and estimate future demand and supply. The optimal lag is selected based on the value that minimizes the mean squared error (MSE) of the RLS predictions.
The results indicate that users of the DAI pool exhibit slower reactions to interest rate changes compared to those in USDC and USDT pools. Specifically, the optimal lag for the DAI demand and supply curves is found to be $\Delta_{\text{DAI, demand}}=26$ and $\Delta_{\text{DAI, supply}}=16$, whereas for USDC and USDT, the optimal lag is $\Delta=1$. Using these optimal shift values, we apply RLS to estimate the evolution of the demand and supply parameters ($\mathtt{a}_{t}^{b},\mathtt{b}_{t}^{b},\mathtt{a}_{t}^{l},\mathtt{b}_{t}^{l}$) over time for each pool.
B.2. Utilization optimization
We now provide the detailed specifications of the experiment used to generate Figure 2(b). The synthetic parameters $\mathtt{a}_{t}^{b},\mathtt{b}_{t}^{b},\mathtt{a}_{t}^{l},\mathtt{b}_{t}^{l}$ evolve via a Gaussian random walk with an update interval of $T=25$, following:
$\mathtt{a}^{b}_{t+T}=\mathtt{a}_{t}^{b}+\epsilon_{t},\quad\epsilon_{t}\sim\mathcal{N}\left(0,\sigma_{\text{trns}}\cdot|\mathtt{a}_{t}^{b}|\right),$
with similar updates for $\mathtt{b}_{t}^{b},\mathtt{a}_{t}^{l},\mathtt{b}_{t}^{l}$. The noise variance is set to $\sigma_{\text{noise}}=1$, while $\sigma_{\text{trans}}$ varies from 0.1 to 1 (1% to 100%). Simulations run for 1000 time steps, averaging results over 50 independent runs. The initial parameter values are:
$\mathtt{a}^{b}_{0}=-10,\quad\mathtt{a}^{b}_{0}=5000,\quad\mathtt{a}^{l}_{0}=500,\quad\mathtt{b}^{l}_{0}=-50.$
Desired utilization is $U^{*}=0.7$ and forgetting factor is $0.8$.
B.3. Robust RLS experiment
We evaluate the robustness of the robust RLS algorithm against adversarial manipulation of demand and supply using two distinct strategies. The metric of interest is the normalized rate deviation, defined as the deviation of the interest rate from the optimal rate that achieves the target utilization $U^{*}=0.7$. The first strategy is an intermittent adversary, which randomly activates in every timeslot with a probability of 0.1 to enforce large noise on demand and supply. When inactive, the noise variance is $\sigma_{\text{noise}}=10$, and when active, it increases to $\sigma_{\text{noise}}=|B_{t}|\cdot\sigma_{\text{attack}}$ for demand and $\sigma_{\text{noise}}=|L_{t}|\cdot\sigma_{\text{attack}}$ for supply, where $\sigma_{\text{attack}}$ varies from 0.2 to 3 (20% to 300% of demand or supply). As shown in Figure 2(a), the normalized rate deviation for plain RLS increases sharply with $\sigma_{\text{attack}}$, while the robust RLS algorithm maintains near-zero deviation, demonstrating resilience to such attacks. The second strategy simulates a persistent and more realistic adversarial borrower, which activates with probability 0.01 and stays active for $T_{\text{attack}}$ steps. The adversary inflates demand using $\mathtt{a}_{t}^{b}r_{t}\cdot\gamma_{\text{adv}}+\mathtt{b}_{t}^{b}$, where $\gamma_{\text{adv}}>1$ exaggerates sensitivity to interest rates, tricking the protocol into lowering rates. Stress tests with $T_{\text{attack}}=100$ and $\gamma_{\text{adv}}\in\{2,5,10,20\}$ show that while plain RLS suffers increasing normalized rate deviation (up to 3.13 for $\gamma_{\text{adv}}=20$), robust RLS keeps it below 0.5, even under extreme attacks
These experiments were conducted using synthetic demand and supply functions, following the experimental setup described in Appendix B.2.
B.4. Forgetting factor and estimation error
The forgetting factor governs the trade-off between adaptivity and precision; therefore, depending on the dynamics of the demand and supply parameters, different forgetting factors should be used. Figure 3 illustrates the mean squared error (MSE) of the parameters $\mathtt{a}_{t}^{b},\mathtt{b}_{t}^{b},\mathtt{a}_{t}^{l},\mathtt{b}_{t}^{l}$ estimated by the RLS-based algorithm. The parameters were generated using a random walk with Gaussian noise, with a frequency of change every $T=50$ steps and $\sigma_{\text{trns}}=0.1$. We run the experiment for $1000$ timeslots and average the results across $50$ independent runs. In this experiment setup, the optimal forgetting factor, which minimizes the overall MSE across all parameters, is approximately $\rho=0.82$.
@@ B.4. Forgetting factor and estimation error
Figure 3. Impact of forgetting factor on the mean square error of the estimated parameters by the RLS-based algorithms.
B.5. Revenue maximizing
@@
Figure 4. Comparison of supply, used as a proxy for revenue, between the RLS-based controller and Aave’s static curves. The parameters evolve according to a random walk with Gaussian noise. The x-axis represents the relative standard deviation of the noise in percentage. The fixed factor is set to $\rho=0.8$.
B.6. Risk controller
Figure 5. ETH volatility over time, with the controller dynamically adjusting the collateral factor. The bottom plot compares the actual liquidations resulting from this adjusted collateral factor with the target expected liquidations, $LT_{t}=0.9$.
Appendix C Proofs
C.1. Proof of theorem 4.1
Proof.
Our proof is applying a slightly modified version of the proof provided in (Canetti and España, 1989). The condition required for convergence of the RLS algorithm is that the input vector of the linear system should be persistently excited. In other words if we want to estimate $\mathbf{\theta}_{t}$, and the linear system is $y_{t}=\mathbf{x}_{t}^{T}\mathbf{\theta}_{t}+e_{t}$ then $\mathbf{x}_{t}$ should hold in:
@@ Appendix C Proofs
The above analysis shows that as long as for any $t$, there is some $N$ such that Algorithm 2 sets at least two distinct interest rates from time $t$ to $t+N$, then the excitation condition of paper (Canetti and España, 1989) is met.
@@
So now we can apply the results of (Canetti and España, 1989) for the convergence of the RLS algorithm to the right $\theta_{t}$. They write the estimation error $\tilde{\theta}_{t}=\theta_{t}-\hat{\theta}_{t}$ as sum of three terms $\tilde{\theta}_{t}\coloneqq\tilde{\theta}^{1}_{t}+\tilde{\theta}^{2}_{t}+\tilde{\theta}^{3}_{t}$. Given that $\theta$ has some major change at time $t_{0}$, the question is at time $t>t_{0}$ what is an upper bound of each of $\tilde{\theta}^{1}_{t},\tilde{\theta}^{2}_{t},\tilde{\theta}^{3}_{t}$? They show that the first and second terms decay exponentially over time i.e., $\|\tilde{\theta}^{1}_{t}+\tilde{\theta}^{2}_{t}\|=\mathcal{O}(\rho^{t})$ whereas the third term $\tilde{\theta}_{t}^{3}$ which is the only term that is a function of the observation noise, is as follows:
$\limsup_{t\to\infty}\mathbb{E}\left(\|\tilde{\theta}_{t}^{3}\|^{2}\right)\leq\nu^{2}\frac{M}{m^{2}}\left(\frac{1-\rho^{N}}{\rho^{N}}\right)^{2}\frac{1}{\ln\left(\frac{1}{\rho}\right)}=\mathcal{O}\left(\left(\frac{1-\rho^{N}}{\rho^{N}}\right)^{2}\frac{\nu^{2}}{\ln{\frac{1}{\rho}}}\right)$
@@
C.2. Proof of theorem 4.2
Proof.
@@
(23) $U$ $=\frac{-\mathtt{a}^{b}\left(R_{\text{slope1}}+R_{\text{slope2}}\left(\frac{U-U^{*}}{1-U^{*}}\right)\right)+\mathtt{b}^{b}}{L}\implies U=\frac{\mathtt{b}^{b}-\mathtt{a}^{b}R_{\text{slope1}}+\frac{\mathtt{a}^{b}R_{\text{slope2}}U^{*}}{1-U^{*}}}{L+\frac{\mathtt{a}^{b}R_{\text{slope2}}}{1-U^{*}}}$
(24) $\implies r=R_{\text{slope1}}+\frac{R_{\text{slope2}}\left(\mathtt{b}^{b}-\mathtt{a}^{b}R_{\text{slope1}}-U^{*}L\right)}{\left(L+\frac{\mathtt{a}^{b}R_{\text{slope2}}}{1-U^{*}}\right)(1-U^{*})}$
@@
$\mathcal{R}_{t}^{\mathcal{P},\nu}(U^{*})=\begin{cases}\left|R_{\text{slope1}}+\frac{R_{\text{slope2}}\left(\mathtt{b}^{b}-\mathtt{a}^{b}R_{\text{slope1}}-U^{*}L\right)}{\left(L+\frac{\mathtt{a}^{b}R_{\text{slope2}}}{1-U^{*}}\right)(1-U^{*})}-\frac{\mathtt{b}^{b}-L\,U^{*}}{\mathtt{a}^{b}}\right|&\text{if }R_{\text{slope1}}<\frac{\mathtt{b}^{b}-L\,U^{*}}{\mathtt{a}^{b}},\\ \\ \left|\frac{\mathtt{b}^{b}R_{\text{slope1}}}{U^{*}\,L+\mathtt{a}^{b}R_{\text{slope1}}}-\frac{\mathtt{b}^{b}-L\,U^{*}}{\mathtt{a}^{b}}\right|&\text{if }R_{\text{slope1}}>\frac{\mathtt{b}^{b}-L\,U^{*}}{\mathtt{a}^{b}}.\end{cases}$
$\mathcal{R}_{t}^{\mathcal{P},\nu}(U^{*})=\begin{cases}\frac{\Delta}{1+\frac{\mathtt{a}^{b}\cdot R_{\text{slope1}}}{L\,U^{*}}}&\text{if }R_{\text{slope1}}<\frac{\mathtt{b}^{b}-L\,U^{*}}{\mathtt{a}^{b}},\\ \\ \frac{\Delta}{1+\frac{\mathtt{a}^{b}\cdot R_{\text{slope2}}}{L(1-U^{*})}}&\text{if }R_{\text{slope1}}>\frac{\mathtt{b}^{b}-L\,U^{*}}{\mathtt{a}^{b}}.\end{cases}$
$\mathcal{R}_{t}^{\mathcal{P},\nu}(U^{*})\geq\frac{\Delta}{1+\max\{\frac{\mathtt{a}^{b}}{L}\}\cdot\max\{\frac{R_{\text{slope1}}}{U^{*}},\frac{R_{\text{slope2}}}{1-U^{*}}\}}$
C.3. Proof of theorem A.1
Proof.
@@
(25) $U_{t}=\frac{\mathtt{b}^{l}+\sqrt{(\mathtt{b}^{l})^{2}-4\mathtt{a}^{l}r_{t}(\mathtt{a}^{b}r_{t}-\mathtt{b}^{b})}}{2\mathtt{a}^{l}r_{t}}$
@@
(26) $r=\frac{\mathtt{b}^{b}+\sqrt{(\mathtt{b}^{b})^{2}+4(\mathtt{a}^{b})^{2}(\mathtt{b}^{l})^{2}}}{2\mathtt{a}^{b}}$
@@
C.4. Proof of Proposition 5.1
Proof.
@@
C.5. Proof of Theorem 5.2
Proof.
@@
(27) $\text{Utility}_{b}^{\text{honest}}\geq$ $-\Bigg{(}\delta_{b}\mathtt{b}^{b}\tau\zeta\frac{r^{b}-r_{\text{min}}}{r_{\text{max}}-r_{\text{min}}}\frac{r^{b}+r_{\text{min}}}{2}+C_{\tau,\zeta}^{r^{b}}+\delta_{b}\mathtt{b}^{b}(T-\tau)r^{b}\Bigg{)}$
If the borrower follows the honest strategy and reports $r^{b}$, during the exploration phase, they will borrow whenever $r\in[r_{\text{min}},r^{b}]$ and pay an average interest rate of $\frac{r_{\text{min}}+r^{b}}{2}$ during those timeslots. There are $\tau\zeta\frac{r^{b}-r_{\text{min}}}{r_{\text{max}}-r_{\text{min}}}$ such timeslots based on the uniform randomization of $\mathcal{C}_{D}$ described in Appendix [A.3]. During the rest of the exploration phase, i.e., when $r>r^{b}$ or during the non-randomized times, we denote the cost of the borrower by $C_{\tau,\zeta}^{r^{b}}$. Finally, during the exploitation phase $T-\tau$, the borrower pays at most rate $r^{b}$ (potentially even lower depending on the rest of the parameters and objective utilization $U^{*}$).
@@
(28) $+C_{\tau,\zeta}^{r^{b}}+\delta_{b}\mathtt{b}^{b}(T-\tau)\bar{r}^{b}\Bigg{)}$
@@
(29) $\text{Utility}_{b}^{\text{adv}}-\text{Utility}_{b}^{\text{honest}}\leq$ $-\delta_{b}\mathtt{b}^{b}\Bigg{(}\frac{\tau\zeta}{2}\frac{(r^{b}-\bar{r}^{b})^{2}}{r_{\text{max}}-r_{\text{min}}}+\big{(}T-\tau\big{)}(\bar{r}^{b}-r^{b})\Bigg{)}$
(30) $\text{Utility}_{b}^{\text{adv}}-\text{Utility}_{b}^{\text{honest}}\geq 0$
(31) $\implies\Bigg{(}\frac{\tau\zeta}{2}\frac{(r^{b}-\bar{r}^{b})^{2}}{r_{\text{max}}-r_{\text{min}}}+\big{(}T-\tau\big{)}(\bar{r}^{b}-r^{b})\Bigg{)}\leq 0$
(32) $\implies 0<r^{b}-\bar{r}^{b}<\frac{2(r_{\text{max}}-r_{\text{min}})}{\zeta}\left(\frac{T}{\tau}-1\right)$
@@
    (33) $\frac{\mathtt{b}^{b}(1-\delta_{b})-\mathtt{a}^{b}\,\bar{r}}{L}<U^{*}\implies\bar{r}>\frac{\mathtt{b}^{b}(1-\delta_{b})-L\,U^{*}}{\mathtt{a}^{b}}$
@@
    (34) $U^{*}-\frac{\mathtt{b}^{b}(1-\delta_{b})-\mathtt{a}^{b}\,\bar{r}}{L}>\frac{\mathtt{b}^{b}-\mathtt{a}^{b}\,\bar{r}}{L}-U^{*}\implies\bar{r}>\frac{\mathtt{b}^{b}(1-\frac{\delta_{b}}{2})-L\,U^{*}}{\mathtt{a}^{b}}$
@@
(35) $\text{Utility}_{l}^{\text{honest}}\geq$ $\Bigg{(}\delta_{l}L\tau\zeta\frac{r_{\text{max}}-r^{l}}{r_{\text{max}}-r_{\text{min}}}\frac{r^{l}+r_{\text{max}}}{2}+C_{\tau,\zeta}^{r^{l}}+\delta_{l}L(T-\tau)r^{l}\Bigg{)}$
@@
(36) $+C_{\tau,\zeta}^{r^{l}}+\delta_{l}L(T-\tau)\bar{r}^{l}\Bigg{)}$
@@
    (37) $\frac{\mathtt{b}^{b}-\mathtt{a}^{b}\,\bar{r}}{L(1-\delta_{l})}>U^{*}\implies\bar{r}<\frac{\mathtt{b}^{b}-L(1-\delta_{l})\,U^{*}}{\mathtt{a}^{b}}$
@@
    (38) $\frac{\mathtt{b}^{b}-\mathtt{a}^{b}\,\bar{r}}{L(1-\delta_{l})}-U^{*}>U^{*}-\frac{\mathtt{b}^{b}-\mathtt{a}^{b}\,\bar{r}}{L}\implies\bar{r}<\frac{\mathtt{b}^{b}-U^{*}\,L\frac{(1-\delta_{l})}{(1-\frac{\delta_{l}}{2})}}{\mathtt{a}^{b}}$
@@
C.6. Proof of Theorem 5.3
Proof.
@@
(39) $\hat{L}_{\text{strtgc}}>0\ \implies r_{o}^{l}\leq\frac{B\alpha}{L(1-\delta_{l})}+\beta$